OpenStudy (anonymous):

3 fair coins are tossed. What's the probability of obtaining at most 2 heads?

OpenStudy (anonymous):

1/8

OpenStudy (anonymous):

how did you get that?

jimthompson5910 (jim_thompson5910):

There are 8 possible ways to toss 3 coins and they are HHH HHT HTH HTT THH THT TTH TTT Of this list, there are 7 ways to get either 1 head or 2 heads So the probability is 7/8

OpenStudy (anonymous):

because when you are doing probability, you need to multiply the chance that it will be heads, which is 1 out of 2 or 1/2. therefore it is 1/2 *1/2 which is 1/4 times 1/2 which is 1/8. this is the probability that all of the answers are heads

jimthompson5910 (jim_thompson5910):

Alternatively, you can compute the probability of getting all tails, which is (1/2)*(1/2)*(1/2) = 1/8 Then subtracting that from 1 to get 1 - 1/8 = 7/8 getting the same answer

OpenStudy (anonymous):

but isn't that just exactly 2 heads and not at most 2 heads?

jimthompson5910 (jim_thompson5910):

oh wait, the probability is still the same, I just mixed up HHH and TTT

OpenStudy (anonymous):

@Jim In the list of possibilities, you have HHH counted as one of the "either 1 or 2". Does that mean that the probability is 6/8?

jimthompson5910 (jim_thompson5910):

I mixed up HHH and TTT on accident I meant to write... There are 8 possible ways to toss 3 coins and they are HHH HHT HTH HTT THH THT TTH TTT Of this list, there are 7 ways to get 0 heads, 1 head or 2 heads (ie "at most 2 heads") Basically, it's everything but HHH. So the probability is 7/8

jimthompson5910 (jim_thompson5910):

So you can see that the probability is still the same.

OpenStudy (anonymous):

oh. TTT is still at most 2 heads. Thank you.

jimthompson5910 (jim_thompson5910):

Yes as 0 heads is less than that upper max limit.

OpenStudy (anonymous):

my bad! i thought that it said all heads! my bad! therefore , the probability is 7/8